Joey is 4 years younger than Sam. If Sam is t years old, how old is Joey? Also, if Steven is twice as old as Joey, how old is he?

Answer by macston(5194)

.
J=Joey's age; S=Sam's age; V=Steven's age
.
J=S-4=t-4
.
V=2J=2(t-4)=2t-8
.